OpenStudy (anonymous):

Please Help Me. I really need it. If you help I'll give you a medal. (: Show an equation and a solution for the problem. A cargo plane left the San Francisco International Airport at noon, traveling east. At the same time, a jet left the Houston Airport traveling west 250 mph faster than the cargo plane. If the airports are 1,403 miles apart and the jet passes the cargo plane in 2.3 hours, what is the speed of each?

OpenStudy (anonymous):

Anna, if u hv not understood the solution given by Randi, let me know and I will explain it in a different manner....

OpenStudy (anonymous):

Can you explain it better.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

let the speed of the cargo plane be = x mph then speed of the jet plane = (x+250) mph Since they leave together and meet after 2.3 hours, each has been travelling for 2.3 hours in opposite direction In 2.3 hours the distance covered by the cargo plane = 2.3x In 2.3 hours the distance covered by jet plane = 2.3(x+250) When they meet, after travelling from opposite directions, their combined distance is equal to the distance between the two airports i.e. 1403 so 2.3x + 2.3(x+250) = 1403 2.3x + 2.3x + 575 = 1403 4.6x = 1403-575 x = 828/4.6 = 180 so speed of cargo plane = x = 180 mph speed of jet plane is 180 + 250 = 430 mph If clear, pls click GOOD ANS button for me ( :})

OpenStudy (anonymous):

???????????????